# Understanding Outer Optimizers in Local SGD: Learning Rates, Momentum, and Acceleration

### v1

@@ Appendix B Guarantees for Local SGD
where $g_{m,r,h}=\nabla f(y_{m,r,h})+n_{m,r,h}$ is the stochastic gradient with noise $n_{m,r,h}$, and $\overline{g}_{m,r,h}=\nabla f(y_{m,r,h})$ is the true gradient. By Assumption 3.2 we have $\mathbb{E}\left[g_{m,r,h}\right]=\overline{g}_{m,r,h}$. After $H$ local steps, the global model update can be equivalently written as $x_{r+1}=x_{r}-\gamma\eta\sum_{h=0}^{H-1}g_{r,h}$ where $g_{r,h}=\frac{1}{M}\sum_{m=1}^{M}g_{m,r,h}$ is the average gradient across workers and $y_{r,h}=\frac{1}{M}\sum_{m=1}^{M}y_{m,r,h}$ is the average model. Note that these two last sequences are virtual sequences and not actually computed. We also define $x_{r,h}=x_{r}-\gamma\eta\sum_{h=0}^{H-1}g_{r,h}$ as an intermediate quantity used in the analysis.
@@ B.1 Algorithm-independent results
Lemma B.1.
[26, Lemma 6] Let $f$ be a convex and $L$-smooth function. Suppose that $\eta\leq\frac{2}{L}$, let $T_{\eta}(x)=x-\eta\nabla f(x)$. Then
@@
By the Baillon-Haddad theorem [1] we have
@@
Lemma B.4.
[21, Lemma 7]. Let $S$ be the set of nonnegative and nondecreasing sequences. Let $y_{1},y_{2},\ldots$ be a sequence in $S$. Let $C_{t}\in\mathcal{F}_{t-1}$ for all $t=1,2,\ldots,T$ and let $X_{t}$ be a martingale difference sequence adapted to $\mathcal{F}_{t}$ such that $\left\lvert X_{t}\right\rvert\leq C_{t}$ with probability $1$ for $t=1,2,\ldots,T$. Then for all $\delta\in(0,1)$ and $\hat{X}_{t}\in\mathcal{F}_{t-1}$ such that $\left\lvert\hat{X}_{t}\right\rvert\leq C_{t}$ with probability $1$, we have that with probability at least $1-\delta-\mathrm{Prob}\left(\exists t\leq T\mid C_{t}>c\right)$ that for all $c>0$
@@
B.2 Non-adaptive guarantee without momentum
@@
Lemma B.6 (Regret against starting point).
@@
Lemma B.7.
(Local client drift bound). Suppose that Assumptions˜3.1 and 3.2 hold. Then in Algorithm GEN-LOC-SGD for all $r$ and $h$, if $\eta\leq\frac{1}{L}$, then
@@
Finally, using the fact that ${\left\lVert T_{\eta}(x)-T_{\eta}(y)\right\rVert}^{2}\leq{\left\lVert x-y\right\rVert}^{2}$ whenever $\eta\leq\frac{2}{L}$ (Lemma B.1) and Assumption 3.2, we get
@@
For the variance term, when $\eta\leq\frac{1}{L}$ we use Lemma˜B.7
@@
We now have two cases. Case 1. If $\gamma\geq 1$, then we have by Lemma˜B.3 and Jensen’s inequality applied to ${\left\lVert\cdot\right\rVert}^{2}$,
@@
Observe that we can write $\max(\gamma,1)=1+(\gamma-1)_{+}$. Dropping the negative term and using Jensen’s inequality gives
@@
B.3 Non-adaptive guarantee with momentum
Proof of Theorem 3.5.
@@ B.3 Non-adaptive guarantee with momentum
Following the same proof as Theorem˜3.3, we can bound (in expectation)
@@
Using Lemma˜B.3 we have
@@ B.4 Acceleration proofs
where $V_{r,h}=\frac{1}{M}\sum_{m=1}^{M}{\left\lVert\nabla f(u_{r,h})-\nabla f(u_{m,r,h})\right\rVert}^{2}\leq\frac{L^{2}}{M}\sum_{m=1}^{M}{\left\lVert u_{r,h}-u_{m,r,h}\right\rVert}^{2}$. Taking unconditional expectation, dropping the ${\left\lVert\nabla f(u_{r,h})\right\rVert}^{2}$ term and using Lemma B.7 we have
@@
Lemma B.9 (Momentum sequence bound).
@@
Taking expectations and using Lemma B.3,
@@
Using Lemma B.8 and Lemma B.9, we have
@@
The proof of Corollary˜3.7 is straightforward by substitution and is omitted for brevity.
@@ B.5 Data-dependent guarantees
Furthermore, we have by the triangle inequality, then our assumption on the noise followed by Lemma B.1 that almost surely
@@
Then by Lemma˜B.4 with $y_{h}=\overline{\mu}_{r,h}$ we have with probability at least $1-\delta$
@@
Observe that by the triangle inequality followed by Lemma˜B.2
@@
Since this holds for any $h$, we have that $\zeta_{2}\leq\eta\sum_{k=0}^{H-1}\parallel g_{r,k}\parallel$, where $\zeta_{2}$ is defined in Lemma˜B.11. Moreover, by Lemma˜B.10 we have that with probability $1-\delta$ and an application of the union bound that for all $r,h$
@@
Therefore by Equation˜53 and Lemma˜B.10
@@
Applying Lemma˜B.4 on $X_{r,h}$ with $y_{r,h}=\bar{\nu}_{r,h}$, $C_{r,h}=\sigma$, $\hat{X}_{r,h}=0$ we have

### v2

@@ Appendix B Guarantees for Local SGD
where $g_{m,r,h}=\nabla f(y_{m,r,h})+n_{m,r,h}$ is the stochastic gradient with noise $n_{m,r,h}$, and $\overline{g}_{m,r,h}=\nabla f(y_{m,r,h})$ is the true gradient. By Assumption 2 we have $\mathbb{E}\left[g_{m,r,h}\right]=\overline{g}_{m,r,h}$. After $H$ local steps, the global model update can be equivalently written as $x_{r+1}=x_{r}-\gamma\eta\sum_{h=0}^{H-1}g_{r,h}$ where $g_{r,h}=\frac{1}{M}\sum_{m=1}^{M}g_{m,r,h}$ is the average gradient across workers and $y_{r,h}=\frac{1}{M}\sum_{m=1}^{M}y_{m,r,h}$ is the average model. Note that these two last sequences are virtual sequences and not actually computed. We also define $x_{r,h}=x_{r}-\gamma\eta\sum_{h=0}^{H-1}g_{r,h}$ as an intermediate quantity used in the analysis. Table˜6 summarizes some of the notation we use throughout this section.
Table 6: Key notation.
Symbol Description Symbol Description
$M$ Number of nodes $x_{r}$ Global iterate at round $r$
$H$ Local steps per round $y_{r,h}$ Averaged local iterate
$R$ Communication rounds $g_{r,h}$ Averaged stochastic gradient
$\eta$ Inner learning rate $L$ Smoothness constant
$\gamma$ Outer learning rate $\sigma^{2}$ Gradient variance bound
$\mu$ Momentum parameter $D$ $\parallel x_{0}-x_{*}\parallel$ initial distance to optimum
@@ B.1 Algorithm-independent results
Lemma 1.
[karimireddy19_scaff, Lemma 6] Let $f$ be a convex and $L$-smooth function. Suppose that $\eta\leq\frac{2}{L}$, let $T_{\eta}(x)=x-\eta\nabla f(x)$. Then
@@
By the Baillon-Haddad theorem [bauschke09_baill_haddad_theor_revis] we have
@@
Lemma 4.
[ivgi23_dog_is_sgds_best_frien, Lemma 7]. Let $S$ be the set of nonnegative and nondecreasing sequences. Let $y_{1},y_{2},\ldots$ be a sequence in $S$. Let $C_{t}\in\mathcal{F}_{t-1}$ for all $t=1,2,\ldots,T$ and let $X_{t}$ be a martingale difference sequence adapted to $\mathcal{F}_{t}$ such that $\left\lvert X_{t}\right\rvert\leq C_{t}$ with probability $1$ for $t=1,2,\ldots,T$. Then for all $\delta\in(0,1)$ and $\hat{X}_{t}\in\mathcal{F}_{t-1}$ such that $\left\lvert\hat{X}_{t}\right\rvert\leq C_{t}$ with probability $1$, we have that with probability at least $1-\delta-\mathrm{Prob}\left(\exists t\leq T\mid C_{t}>c\right)$ that for all $c>0$
@@
B.2 Convergence guarantees without momentum
@@
Lemma 6 (Regret against starting point).
@@
Lemma 7.
(Local client drift bound). Suppose that Assumptions˜1 and 2 hold. Then in Algorithm GEN-LOC-SGD for all $r$ and $h$, if $\eta\leq\frac{1}{L}$, then
@@
Finally, using the fact that ${\left\lVert T_{\eta}(x)-T_{\eta}(y)\right\rVert}^{2}\leq{\left\lVert x-y\right\rVert}^{2}$ whenever $\eta\leq\frac{2}{L}$ (Lemma 1) and Assumption 2, we get
@@
For the variance term, when $\eta\leq\frac{1}{L}$ we use Lemma˜7
@@
We now have two cases. Case 1. If $\gamma\geq 1$, then we have by Lemma˜3 and Jensen’s inequality applied to ${\left\lVert\cdot\right\rVert}^{2}$,
@@
Dropping the negative term and using Jensen’s inequality gives
@@
B.3 Convergence guarantees with momentum
Proof of Theorem 2.
@@ B.3 Convergence guarantees with momentum
Following the same proof as Theorem˜1, we can bound (in expectation)
@@
Using Lemma˜3 we have
@@ B.4 Acceleration proofs
where $V_{r,h}=\frac{1}{M}\sum_{m=1}^{M}{\left\lVert\nabla f(u_{r,h})-\nabla f(u_{m,r,h})\right\rVert}^{2}\leq\frac{L^{2}}{M}\sum_{m=1}^{M}{\left\lVert u_{r,h}-u_{m,r,h}\right\rVert}^{2}$. Taking unconditional expectation, dropping the ${\left\lVert\nabla f(u_{r,h})\right\rVert}^{2}$ term and using Lemma 7 we have
@@
Lemma 9 (Momentum sequence bound).
@@
Taking expectations and using Lemma 3,
@@
Using Lemma 8 and Lemma 9, we have
@@
The proof of Corollary˜1 is straightforward by substitution and is omitted for brevity.
@@ B.5 Data-dependent guarantees
Furthermore, we have by the triangle inequality, then our assumption on the noise followed by Lemma 1 that almost surely
@@
Then by Lemma˜4 with $y_{h}=\overline{\mu}_{r,h}$ we have with probability at least $1-\delta$
@@
Observe that by the triangle inequality followed by Lemma˜2
@@
Since this holds for any $h$, we have that $\zeta_{2}\leq\eta\sum_{k=0}^{H-1}\parallel g_{r,k}\parallel$, where $\zeta_{2}$ is defined in Lemma˜11. Moreover, by Lemma˜10 we have that with probability $1-\delta$ and an application of the union bound that for all $r,h$
@@
Therefore by Equation˜53 and Lemma˜10
@@
Applying Lemma˜4 on $X_{r,h}$ with $y_{r,h}=\bar{\nu}_{r,h}$, $C_{r,h}=\sigma$, $\hat{X}_{r,h}=0$ we have